OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point:2, -6 ; Slope: -3/4

pooja195 (pooja195):

\[y-y1=m(x-x1)\] plug in the numbers like we did before

OpenStudy (anonymous):

i think it is y+6= -3/4(x+2) ??

pooja195 (pooja195):

yep :)

OpenStudy (anonymous):

identify the point (x1, y1) from the equation: y + 8 = 3(x – 2) Is it -2, 8?

pooja195 (pooja195):

what are the coordinates?

OpenStudy (anonymous):

it doesnt show me any. all it asks is 4. Identify the point (x1, y1) from the equation: y + 8 = 3(x – 2) (Points : 1) (–2, 8) (8, –2) (–8,2) (2, –8)